OpenStudy (anonymous):

The physics department of a college has 9 male professors, 8 female professors, 10 male teaching assistants, and 13 female teaching assistants. If a person is selected at random from the group, find the probability that the selected person is a teaching assistant or a female.

OpenStudy (anonymous):

Number who are teaching assistants and/or female 10 + 13 + 8 ------------------------------------------ = ------------------ Total Number of People 9 + 8 + 10 + 13

OpenStudy (mathmale):

BG: Would you mind typing out the formula used to calculate P(A or B), when P(A), P(B) and P(A AND B) are given?

OpenStudy (anonymous):

In my opinion, students should not automatically think "Which formula do I need for this?" I tend not to explain things that way. For many it makes intuitive sense to do these applying the concept. Number of things we want ---------------------- Total number of things we're choosing from

OpenStudy (mathmale):

My result is dramatically different from yours: \[\frac{ 23 }{ 40 }+\frac{ 21 }{ 40 }-\frac{ 13 }{ 40 }=\frac{ ? }{ 40 }\]

OpenStudy (mathmale):

So long as we arrive at the same answers, I'm not so concerned about using a formula versus using an intuitive approach.

OpenStudy (mathmale):

BG: I first set up a contingency table, which made it easy to determine P(TA), P(F) and P(TA and F).

OpenStudy (mathmale):

@Elaina928: May we hear from you, please? It's important that you be involved in finding the solution to what is, after all, your math problem.